OpenStudy (anonymous):

x - 3y if x = 3 and y = -2.

OpenStudy (anonymous):

put x=3 and y=-2, x-3y=3-3(-2)=3+6=...

OpenStudy (unklerhaukus):

\[x - 3y\] , \[3-3\times-2\]

OpenStudy (anonymous):

3 - (-6)

OpenStudy (anonymous):

9